Consider the following preemptive variant of the Load Balancing problem: There are m identical machines, and n jobs with processing times t1, t2,...., tn respectively. Each machine can process at most one job at a time. Each job can run on more than one machines but must run on at most one machine at any time. Let T = max { max (1<=j<=n) tj , 1/ m "summation" (1<=j<=n) tj }

1
Expert's answer
2021-05-02T08:35:58-0400

Given:

Consider the following preemptive variant of the Load Balancing problem: There are m identical machines, and n jobs with processing times "t_1,t_2,...,t_n" respectively. Each machine can process at most one job at a time. Each job can run on more than one machines but must run on at most one machine at any time. Let"T = \max \left\{
 \max\limits_{(1<=j<=n)} t_j , \frac{1}{ m} \sum\limits_{(1<=j<=n)} t_j \right\}"

Solving:

"\forall i" Plot in line section length "t_i" ,that end i section=start i+1 section



in the result we give section with length= "\sum\limits_{(1<=j<=n)} t_j".

this section divided by section length T.

Because T*m "\geq\sum\limits_{(1<=j<=n)} t_j" we give m or less section.



 transform this graphic for machines

"p_i -process(job)\ with\ index\ i"

"m_i-machine\ with\ index\ i"



Because "\forall j T\geq t_j ,p_j" -run on 1 or 2 machines . If "p_j" run on 1 machines , we didn't problem ,"p_j" - finished and  run on at most one machine at any time.if "p_j" run on 2 mashines "p_j" run on at most one machine at any time, because if exsist time ,that "p_j"  run on most than one machine in this time ,therefore "t_j>T" ,but "t_j\leq T"